OpenStudy (anonymous):

I am mucho of the confused. How do inverse functions? Given f(x) = the quantity of 5x plus 1, divided by 2 solve for f^-1(8).

OpenStudy (jdoe0001):

\(\bf f(x)={\color{red}{ y}}=\cfrac{5{\color{blue}{ x}}+1}{2}\qquad inverse\implies {\color{blue}{ x}}=\cfrac{5{\color{red}{ y}}+1}{2}\Leftarrow f^{-1}(x)\) as you can see, to get the inverse relation, we simply swap about the variables then solve for "y"

OpenStudy (anonymous):

How do you solve for y like that then? I mean like... Do you end up multiplying by 2, to then get 10y+2 and then -2 from each side.. Or like, what...?

OpenStudy (jdoe0001):

well, yes you do end up multiplying by 2 first then subtract 1 from both sides then divide by 5

OpenStudy (anonymous):

2(5y+1) 10y+2??

OpenStudy (jdoe0001):

\(\bf x=\cfrac{5y+1}{2}\implies 2x=5y+1\implies 2x-1=5y\\ \quad \\\implies \cfrac{2x-1}{5}=y\Leftarrow f^{-1}(x)\)

OpenStudy (jdoe0001):

\(\bf f^{-1}({\color{red}{ 8}})=\cfrac{2({\color{red}{ 8}})-1}{5}\)

OpenStudy (anonymous):

OOoooohhhhh, I understand it now-- Thank you very much!!

OpenStudy (jdoe0001):

yw